OpenStudy (mendicant_bias):

Problem regarding magnetic field strength, prompt posted below shortly.

OpenStudy (mendicant_bias):

A wire with mass per unit length 75 grams per meter runs horizontally at right angles to a horizontal magnetic field. A 6.2 Amp current in the wire results in its being suspended against gravity. What's the magnetic field strength?

OpenStudy (mendicant_bias):

The first thing I'd be thinking of doing would be finding the total mass of the wire based on its length so I can figure out F=mg, where mg is equal to the force suspending it against gravity, but I *can't* do that, because I'm not given an overall length, just a mass per unit length.

OpenStudy (mendicant_bias):

@Pompeii00

OpenStudy (anonymous):

haha! they want you to be clever here!

OpenStudy (anonymous):

Can you set up the entire equality for the magnetic force versus the weight of the wire in the most general way possible?

OpenStudy (anonymous):

then maybe you'll see how you can get the answer

OpenStudy (mendicant_bias):

Alright, one sec, let me think about it.

OpenStudy (mendicant_bias):

\[qV\times B - ma=0\]

OpenStudy (anonymous):

Okay, now try to write out the cross product as the magnitude of the cross product.

OpenStudy (mendicant_bias):

\[F_{B}-F_{g}=0\]

OpenStudy (kainui):

Yeah, if they give you the mass per unit length then you probably need to find something like the magentic field strength per unit length to counteract it. Or just go through it with a mass and hope that later it divides out.

OpenStudy (mendicant_bias):

Alright, will do.

OpenStudy (mendicant_bias):

\[qVB \sin(\theta)=qVB; \ \ \ qVB-mg=0.\]

OpenStudy (anonymous):

Very nice!

OpenStudy (anonymous):

Now move the mg to the right side, and then go ahead and rewrite the left in terms of different variables

OpenStudy (mendicant_bias):

Alright, and now just moving mg over to the other side and dividing by qV.

OpenStudy (mendicant_bias):

Yep, heh

OpenStudy (mendicant_bias):

\[B=\frac{mg}{qV}\]

OpenStudy (anonymous):

okay... now qV can actually be rewritten as different quantities!

OpenStudy (mendicant_bias):

Their answer, though, is in the form \[\frac{IlB}{mg}\]

OpenStudy (mendicant_bias):

Which seems like a wholly different situation to me but I guess is the same thing/I dunno? Trying to figure that out

OpenStudy (mendicant_bias):

Whoops, reciprocal of what I just wrote, sorry

OpenStudy (mendicant_bias):

\[B=\frac{mg}{Il}\]Are these the equivalent substitutions?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

because you can rewrite v = L/t then move the t under the q, q/t = I so qv = I L

OpenStudy (mendicant_bias):

\[V = IR; \ qV = qIR. \] \[v=L/t, \]Lost you on that last formula there. Where did L/t come from?

OpenStudy (anonymous):

velocity is length divided by time (meters/second)

OpenStudy (mendicant_bias):

Oh, I thought you meant Inductance for a second, capital L, and got all confused.

OpenStudy (anonymous):

haha sorry, i was trying to distinguish... cuz I and I are the same, even though ones 'i' and one's L haha

OpenStudy (anonymous):

but yeah, then you have m/L, which is your mass density

OpenStudy (anonymous):

so you have all you need

OpenStudy (mendicant_bias):

Alright, yeah, this makes sense. Thank you. Needed to look back on the substitutions for a moment to make sure I got everything.